OpenStudy (anonymous):

determine the vertex and the axis of symmetry for the function Y=-x^2+2x+1

OpenStudy (anonymous):

\[Y=-x^2+2x+1\]

OpenStudy (anonymous):

Can Somebody help me please D:

OpenStudy (anonymous):

same magic formula as before. use \[\frac{-b}{2a}\] for the first coordinate of the vertex. second coordinate is what you get when you plug it in

OpenStudy (anonymous):

in this case a = -1, b = 2 so \[-\frac{b}{2a}=-\frac{2}{2\times(-1)}=1\]

OpenStudy (anonymous):

plug in 1 for x and you win. i get -1+2+1=2 so vertex is (1,2)